IPP Progress Report for the week YYYY.MM.DD - YYYY.MM.DD

(Up to IPP Progress Reports)

Eugene Magnier

Serge Chastel


  • "files monitoring"-related recovery after MHPCC blackout. Many tables on the storage nodes have been corrupted
  • ippdb00 replacement disk plan


  • Regularly kicked the IPP to have exposures (mainly OSS) published


  • mops14 python and condor install: has its own pool
  • Study of iasc detections in smf
  • Meetings with Larry and Denver

Heather Flewelling

  • so i've been playing with skyprobe recently, and the result of which is the following;
    • we suck at stacks on skyprobe. I took 200 or so exposures and stacked in IRAF. It looked much better http://svn.pan-starrs.ifa.hawaii.edu/trac/ipp/wiki/HeathersStackingAdventures
    • we suck at stacks on gpc1. I grabbed some gpc1 warps for SAS2, and stacked (same ones as that went into the stacks) using iraf, there is more coverage with iraf than other. Ken took over my computer and did further gpc1/iraf tests
    • astrometry mysteriously got better recently (?) - skyprobe's astrometry is magically better since the last time I did that in september. No idea why, the recipes are the same as before
    • there were many changes in the gpc1 database since sept, not all of them are documented, which made it hard to fix skyprobe
    • there were changes in ipp which made skyprobe (and non gpc1 stuff) unprocessable. I've fixed what I can, but only for camera / warp /stack stage.
    • there were some changes needed to ingest skyprobe in to a dvodb - not that I trust the astrometry in it because it was the processing from sept.
    • skyprobe debugging of datastore: the summit datastore is funny and reports changing bytes/md5sum for some skyprobe images (generally, while they are copying to the summit datastore) - this breaks summitcopy for us.
    • and... discovery - when the skyprobe is warm the astrometry is crappy.
  • various addstar related stuff has been in process
    • slice00 and slice02 are addstarring
    • slice 12 and slice 16 are merged (slice 14 is next)
    • slice 18 is being delstarred - this takes forever.
  • various czar related tasks.

Mark Huber

  • czar monday, tuesday, extensive monitoring all week/weekend to push MD08 through... misc liaison work continued
  • MD08 refstack week through weekend with system problems. MD08.pv1 started but put on hold while system having problems
  • WW diffim problems -- scanning logs for any good uses for dual conv, waiting for specific position+dates of any examples of missed events due to diffims

Bill Sweeney

Chris Waters

  • Solved stacking weight/detection limit issue. The root cause was that the PSF matching code was not always scaling all input exposures to match. This would have been fine, but we were also not using an average of all inputs to generate the target PSFs. This led to exposures being weighted incorrectly, resulting in the excess detection issue. This is now resolved, and I'm moving on to tuning the FWHM rejection to allow more inputs to be retained.

Chris Waters